OpenStudy (anonymous):

Simple Interest I=Prt Principal($) - 500 Rate(%) - Unknown Time(Yrs) - 5 Interest Earned($) -62.50 Solve for the unknown. Please show work and explain.

Parth (parthkohli):

Simply use the formula to get this:\[62.50 = 500 \times \dfrac{r}{100}\times 5\](Remember that percentage is just something over 100.)

Parth (parthkohli):

Hey, I just wrote an equation, and it's pretty easy. So above is the equation... solve for \(r\), maybe?

OpenStudy (anonymous):

Okay, doing so right now.